# Neither Valid nor Reliable? Investigating the Use of LLMs as Judges

### v1

@@ 2 Large Language Models as Judges
Li et al. (2024b) formalize the evaluation process of LLJs as follow:
$(\mathcal{Y},\mathcal{E},\mathcal{F})=E(\mathcal{T},\mathcal{C},\mathcal{X},\mathcal{R})$ (1)
where $E$ is what they call the evaluation function, taking the evaluation type $\mathcal{T}$, evaluation criteria $\mathcal{C}$, evaluation item $\mathcal{X}$ and optional references $\mathcal{R}$ as inputs. The evaluation function $E$ can correspond to a single LLJ, multiple LLJs (usually referred to as LLM Juries) or a LLJ with a human in the loop. The evaluation type $\mathcal{T}$ is typically pointwise (one item at the time), pairwise or listwise (relative comparison), the evaluation criteria $\mathcal{C}$ is specific to the task at hand and refers to linguistic quality, content accuracy or task-specific metrics. Finally, the evaluation can be reference-based $\mathcal{R}$ or reference-free. The evaluation function can produce three outputs: the evaluation result $\mathcal{Y}$—that can take different forms: e.g, a score, ranking, label or qualitative assessment; the explanation $\mathcal{E}$ of the reasoning that led to the assessment; and the feedback $\mathcal{F}$ consisting of suggestions to improve the input. This formalization accounts for the high variety of LLJ paradigms, we refer to Li et al. (2024b)’s work on the topic for a comprehensive survey on LLJs.
Although early work on LLJs introduced them as an alternative to traditional NLG evaluation metrics (Wang et al., 2023; Liu et al., 2023; Kocmi and Federmann, 2023), their use has since expanded to different applications. Li et al. (2024b) identify three main functionalities: (1) performance evaluation, (2) model enhancement, and (3) data construction.
Performance evaluation refers to the conventional use of LLJs to assess model outputs or overall performance. Model enhancement captures the use of LLJs to improve models, either through reward modeling or by providing feedback throughout the training process. Lastly, data construction involves leveraging LLJs for tasks such as data annotation or data generation. In this work, we examine the use of LLJs along three use cases to ground our analysis, each corresponding to one of these functionalities: text summarization, safety alignment, and data annotation.
@@ 4.2 Assumption 2: LLMs as Capable Evaluators
Example: LLJs as Annotators. Because of LLMs ability to produce human-like text often indistinguishable from human-written text (Clark et al., 2021) and even perceived as higher quality (Zhang and Gosline, 2023; Silva et al., 2024; Parshakov et al., 2025), they are often seen as a great alternative for automated annotation, despite some researchers still disagreeing on their proficiency on the task (Koo et al., 2024; Nasution and Onan, 2024). Interestingly, LLJs have been proposed as substitutes for humans in highly subjective and contested tasks, such as hate speech detection (Huang et al., 2023) and political affiliation classification (Törnberg, 2024). The contested nature of these tasks makes it challenging to establish the content validity of LLJs as annotators. These constructs are highly subjective (Hettiachchi et al., 2023; Goyal et al., 2022), and relying on LLJs—who tend to yield higher inter-annotator agreement and present their own inherent biases—risks overlooking the valuable diversity found in human disagreement, which is especially important in these contexts (Fleisig et al., 2023).
@@ 5 The Path Forward
First, despite the wide range of applications for LLJs, there has been little adaptation in how they are deployed or in how evaluations are designed across different tasks and domains—an oversight that can lead to harmful consequences. For instance, while using LLJs to scale red-teaming efforts can enhance the breadth of evaluation, applying the same approach within a safety mitigation pipeline can result in superficial safety behaviors. In such contexts, more targeted and domain-specific mitigation strategies can be more effective than a trial-and-error approach that does not account for the sociotechnical aspect of safety (Chehbouni et al., 2025). Evaluating the role of LLJs as evaluators necessitates a comprehensive approach that considers several critical dimensions, including the task at hand, the field of application, and the goal of the evaluation, among other things.

### v2

@@ 2 Large Language Models as Judges
Li et al. (2024b) formalize the evaluation process of LLJs as an evaluation function that takes various inputs. The evaluation function can be a single LLJ, multiple LLJs (usually referred to as Juries), or a LLJ with a human in the loop. This function takes as input: an evaluation type: pointwise (one item at the time), pairwise or listwise; an evaluation criteria: specific to the task at hand and refers to linguistic quality, content accuracy or task-specific metrics; an evaluation item; an optional reference: the evaluation can be reference-based or reference-free. This evaluation function can produce three outputs: the evaluation result: e.g, a score, ranking, label or qualitative assessment; the explanation: an explanation of the reasoning that led to the assessment; the feedback: suggestions to improve the input. This formalization accounts for the high variety of LLJ paradigms, we refer to Li et al. (2024b)’s work on the topic for a comprehensive survey on LLJs. Furthermore, while different types of LLJs may exhibit different biases, our analysis in this paper deliberately adopts a higher-level perspective. This enables us to provide a broader perspective on their use and introduction in the field, rather than limiting the discussion to biases tied to specific evaluation types.
Although early work on LLJs introduced them as an alternative to traditional NLG evaluation metrics (Wang et al., 2023; Liu et al., 2023; Kocmi and Federmann, 2023), their use has since expanded to different applications. Li et al. (2024b) identify three main functionalities: (1) performance evaluation, (2) model enhancement, and (3) data construction. Performance evaluation refers to the conventional use of LLJs to assess model outputs or overall performance. Model enhancement captures the use of LLJs to improve models, either through reward modeling or by providing feedback throughout the training process. Lastly, data construction involves leveraging LLJs for tasks such as data annotation or data generation. In this work, we examine the use of LLJs along three use cases to ground our analysis, each corresponding to one of these functionalities: text summarization, safety alignment, and data annotation.
@@ 4.2 Assumption 2: LLMs as Capable Evaluators
Example: LLJs as Annotators. Because of LLMs ability to produce human-like text often indistinguishable from human-written text (Clark et al., 2021) and even perceived as higher quality (Zhang and Gosline, 2023; Silva et al., 2024; Parshakov et al., 2025), they are often seen as a great alternative for automated annotation, despite some researchers still disagreeing on their proficiency on the task (Koo et al., 2024; Nasution and Onan, 2024). Interestingly, LLJs have been proposed as substitutes for humans in highly subjective and contested tasks, such as hate speech detection (Huang et al., 2023) and political affiliation classification (Törnberg, 2024). The contested nature of these tasks makes it challenging to establish the content validity of LLJs as annotators. These constructs are highly subjective (Hettiachchi et al., 2023; Goyal et al., 2022), and relying on LLJs—who tend to yield higher inter-annotator agreement and present their own inherent biases—risks overlooking the valuable diversity found in human disagreement, which is especially important in these contexts. While disagreement is often treated as a marker of poor annotation quality, studies have shown it can instead provide important insights for subjective tasks (Fleisig et al., 2023; Khurana et al., 2024).
@@ 5 The Path Forward
First, despite the wide range of applications for LLJs, there has been little adaptation in how they are deployed or in how evaluations are designed across different tasks and domains—an oversight that can lead to harmful consequences. For instance, while using LLJs to scale red-teaming efforts can enhance the breadth of evaluation, applying the same approach within a safety mitigation pipeline can result in superficial safety behaviors. In such contexts, more targeted and domain-specific mitigation strategies can be more effective than a trial-and-error approach that does not account for the sociotechnical aspect of safety (Chehbouni et al., 2025). Evaluating the role of LLJs as evaluators necessitates a comprehensive approach that considers several critical dimensions, including the task at hand, the goal of the evaluation, and the type of LLJ being used, among other things.